OpenStudy (anonymous):

Radical equations, simple but im stuck with the signs! do you get rid of the negative when you square??

OpenStudy (anonymous):

\[\sqrt{6x+1}+3=0\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

The solution is complex

OpenStudy (anonymous):

i got 10/6

OpenStudy (anonymous):

something is wrong then.

OpenStudy (anonymous):

:D thanks, can someone solve it so i can see if i got the right answer

OpenStudy (anonymous):

\[\sqrt{11} + 3 \neq 0\]

OpenStudy (anonymous):

i think its 9/6? does that sound right

OpenStudy (anonymous):

There is no real number that you can plug that has a negative square root, you need that \[\sqrt{somenumber} = -3\]

OpenStudy (anonymous):

are you agree?

OpenStudy (anonymous):

I'm sorry but I'm confused

OpenStudy (anonymous):

Well the step in wich you square it's wrong, for example: \[2\neq -2\] but \[(2)^2 = (-2)^2\]

OpenStudy (saifoo.khan):

Answer is NO SOLUTION.

OpenStudy (anonymous):

Are you 100% sure?

OpenStudy (saifoo.khan):

200%

OpenStudy (anonymous):

ok thank you, brb

OpenStudy (anonymous):

There is a solution, but its a complex solution.

OpenStudy (saifoo.khan):

Prove it @victor.

OpenStudy (anonymous):

what do you want as a proof?

OpenStudy (anonymous):

if you multiply a negative by a negative its positive....do you rmbr that stuff?